OpenStudy (anonymous):

The radioactive isotope carbon-14 has a half-life of 5730 years. The decay of carbon is?

OpenStudy (anonymous):

i put (ln(.5))/(5730) and it says its wrong....

OpenStudy (unklerhaukus):

\[_6^{14}\text C(t)=~_6^{14}\text C_0\cdot2^{{-t}/{5730\text {yr}}}\]

OpenStudy (anonymous):

where did the 2 come from?

OpenStudy (unklerhaukus):

half life

OpenStudy (anonymous):

i thought the equation for half life is 2=(ln(.5))/(-lambda)

OpenStudy (anonymous):

i mean t=(ln(.5))/(-lambda)

OpenStudy (anonymous):

is this equation right?

OpenStudy (unklerhaukus):

\[t_{1/2}=\frac{\ln(2)}{\lambda}=\frac{-\ln(2)}{-\lambda}=\frac{\ln(2^{-1})}{-\lambda}=\frac{\ln(0.5)}{-\lambda}\]

OpenStudy (anonymous):

so it is correct.....so that means 5730=(ln(.5))/(-lambda)

OpenStudy (anonymous):

i then multiplied the lambda on both sides to get (-lambda)*(ln(.5))=5730

OpenStudy (anonymous):

i mean (5730)*(-lambda)

OpenStudy (anonymous):

then divided by 5730 and got lamda = (ln(.5))/(5730)

OpenStudy (anonymous):

somebody please?

OpenStudy (unklerhaukus):

\[t_{1/2}=\frac{\ln(2)}{\lambda}\] \[{\lambda}=\frac{\ln(2)}{t_{1/2}}\approx\frac{0.693}{5730}=\] about one in the thousand atoms decay every second

OpenStudy (unklerhaukus):

*about one in ten thousand atoms decay every second

OpenStudy (anonymous):

so its ln(2) not ln(.5)?

OpenStudy (anonymous):

can you help me with something another one please?

OpenStudy (unklerhaukus):

i think you just dropped a negative sign somewhere

OpenStudy (unklerhaukus):

i can try

OpenStudy (anonymous):

the derivative of P=.6P(t) and P(0)=1.....How large is the population after 2 years? and How quickly is the population growing after 2 years?

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

for the first one I put e^1.2.,...but wrong